# From Human Hands to Robot Arms: Manipulation Skills Transfer via Trajectory Alignment

### v1

@@ 4 Experiments
Data Collection Efficiency Analysis. As presented in Table 1, a summary of our data collection effort reveals a significant disparity in collection efficiency between human hand and robot teleoperation. For each demonstration, we measure both Collect Time (the execution duration of the task itself) and Reset Time (the period needed to restore the robot, human and objects for the next trial). Overall, Collect Time from a human expert is roughly 3.5 times faster per demonstration than robot teleoperation (an average of 5.09s vs. 17.98s).
@@
Table 1: Statistics of the collected demonstration data. We report the number of demonstrations (# Demos), time cost in minutes, and the data collection efficiency in demonstrations per second (#/s). Efficiency is reported for the collection phase, reset phase, and the total process.
@@
Task Progress (TP): For the more complex, long-horizon tasks (stack the rings on the pillar and stack the paper cups), a simple binary success metric can be too sparse. We therefore measure task progress to provide a finer-grained evaluation. Each task is decomposed into 8 key waypoints, and the policy earns one point for each waypoint it successfully completes. The final score is the average number of completed waypoints across all trials, normalized by the total of 8 waypoints. The detailed waypoint definitions are provided in Appendix A.4.2.
Evaluation Protocol in Real-World Tasks. Each task was evaluated over 50 trials with varied initial object configurations (position, orientation) and the introduction of distractors to test for generalization. Specific randomization details for each task are provided in Appendix A.4.1.
4.1 Performance Analysis
@@
Table 2: Performance comparison. Success Rate (%) for Short-Horizon (SH) tasks and Task Progress (%) for Long-Horizon (LH) tasks are reported. Baseline represents the pre-trained $\pi_{0}$ with task-specific finetuning.
Baseline. We establish a strong baseline using the $\pi_{0}$ architecture (Black et al., ), initialized with its publicly available pre-trained weights. In line with standard VLA evaluation (Cadene et al., 2024), we then fine-tune this model for each downstream task using our collected robot demonstrations. To ensure a fair and direct comparison, the amount of robot data used for fine-tuning the baseline is identical to the robot data subset used for training our full Traj2Action model. This setup provides a rigorous performance benchmark to precisely measure the benefits introduced by our trajectory-guided architecture and the integration of human data.
Contribution of Trajectory Expert. We conduct a two-part ablation study to dissect the contributions of trajectory expert. First, we analyze the impact of our architectural modification. To do this, we evaluate a version of our framework that includes the trajectory expert to baseline but is trained solely on robot data. As shown in Table 2, simply adding the trajectory expert improves performance. The benefit is most pronounced on the long-horizon SPC task, where this architectural change alone boosts the Task Progress score from 37.75% to 54.25% (+16.50%). We attribute this to the explicit decomposition of the policy: the trajectory expert acts as a high-level planner by generating a coarse spatio-temporal plan. This plan serves as a robust prior that simplifies the problem for the action expert, allowing it to focus on refining the motion into precise, low-level actions.
@@
Contribution of Human Data. We then analyze the impact of incorporating human demonstrations for robot learning. As shown in Table 2, integrating human data provides a further, substantial performance boost across all tasks. Compared to the baseline, our full Traj2Action model achieves significant gains: +28% on the pick up the water bottle task (reaching 76.00%), +26.00% on the pick up the tomato and put it in the tray task (reaching 76.00%), +23.50% on the stack the paper cups task (reaching 61.25%), and +21.00% on the stack the rings on the pillar task (reaching 44.75%). This demonstrates that long-horizon tasks, in particular, benefit from the diversity of human data. This is because human demonstrators can intuitively execute a wider range of motions—often involving agile re-orientations or angles that are awkward and difficult to perform via teleoperation. By supplementing the robot dataset with these more varied human demonstrations, the trajectory expert learns to predict more robust and generalizable plans. Consequently, the action expert’s performance is boosted. This effect is also qualitatively illustrated in Figure 4. In the long-horizon stack the rings on the pillar task, the baseline’s poor long-term plan causes it to get trapped. In contrast, our policy predicts a coherent trajectory and executes the task successfully. Furthermore, our policy demonstrates superior robustness: after an initial error in the pick up the tomato and put it in the tray task, it successfully re-plans and recovers, unlike the baseline which enters a non-productive loop.
Comparative video demos of the baseline and our Traj2Action method are provided in the supplementary material. Furthermore, additional videos featuring detailed analyses are available on our website https://anonymous.4open.science/w/Traj2Action-4A45/.
Figure 4: Visual comparison of trajectory and action predicition of short- and long-horizon tasks pick up the tomato and put it in the tray (top) and stack the rings on the pillar (bottom), respectively.
4.2 Ablation Studies
@@
Strategy
Robot Data
(#/min)
Human Data
(#/min)
Total Data
Time (min)
Performance
SR(%)
Baseline
408/202.29
202.29
50
+ Trajectory Expert
+ Robot Data-Only
408/202.29
0/0
202.29
60
+ Trajectory Expert
+ Human & Robot Data
270/133.70
120/14.25
147.95
58
240/28.61
162.31
60
635/75.39
209.09
62
Figure 6: Performance comparison for the pick up the tomato and put it in the tray task under different data collection strategies. The Robot and Human Data columns show the number of demonstrations collected and the total time required in minutes.
@@
Human Data act as an Effective Substitute for Robot Data. To investigate whether labor-saving collected human data can replace time-consuming robot data without degrading policy training performance, we conduct an experiment with three configurations as presented in Table 7.
The results show that human data can be leveraged to train robot policy efficiently. By leveraging just 240 human demonstrations and less robot data, our policy achieves an identical 60% performance compared with model trained solely by robot data, but with a 20% data collection time reduction.
@@
Impact of Trajectory Sampling Frequency. A critical challenge in unifying human and robot data is the inherent mismatch in their kinematic speeds; human movements are typically much faster than robot motions. We hypothesized that to sample trajectory data from human and robot with a consistent speed profile (i.e., where the spatial distance between consecutive points is similar), it is necessary to sample the faster human motion at a higher frequency than the slower robot motion.
@@ 5 Conclusion
6 Reproducibility statement
To ensure the reproducibility of our research, we provide comprehensive details on our methodology, code, and experimental setup.
Code
All code used for data processing, model training, and evaluation is publicly available in our anonymized repository at https://anonymous.4open.science/r/Traj2Action-4A45/. The repository includes instructions for setting up the environment and running the experiments.
Data Collection
A detailed description of the hardware and software setup for both the human hand motion capture and the robot data collection systems is provided in Appendix A.2.
Training
We provide exhaustive details on the training procedures, including hyperparameters, model architectures, and training schedules, in Section 3 and Appendix A.5.
Evaluation
The standardized protocol used for evaluating our policy across all four manipulation tasks, including the definition of metrics and the setup for initial state randomization, is detailed in Section 1 and Appendix A.4.
@@ Appendix A Appendix
A.1 Statement on the Use of Large Language Models (LLMs)
In accordance with the ICLR 2026 policy, we disclose the use of a Large Language Model (LLM) in the preparation of this manuscript. The LLM was employed exclusively as a tool for language polishing, including grammar correction, style improvement, and enhancement of readability. The core scientific contributions, including all research ideas, methodologies, experimental designs, data analysis, and conclusions, were conceived and executed entirely by the human authors. The LLM did not contribute to any substantive aspect of the research. The authors assume full responsibility for all content presented in this paper.
A.2 Data Collection System Details
@@
A.2.1 Human Hand Motion Capture System
@@
Hardware Setup. The system is built around four synchronized cameras operating at 30 Hz (as shown in Figure 2). Three of these are static industrial cameras mounted on a rigid frame surrounding the data collection workspace, providing top-down, side-left, and side-right perspectives of the user’s hand. This arrangement is critical for minimizing self-occlusion. The fourth is a lightweight camera mounted on the back of the user’s hand, providing an ego-centric view analogous to the wrist camera on our robot.
@@
  • 2D Keypoint Detection: We leverage the robust Google MediaPipe (Zhang et al., 2020) Hand Landmarker to process each video stream independently. For each hand detected in an image, the model outputs 21 2D landmarks ($\mathbf{u},\mathbf{v}$) in pixel coordinates. These landmarks serve as the 2D evidence for our 3D reconstruction.
  • Model-Based 3D Reconstruction: We fit the MANO (Romero et al., 2017) model to these multi-view 2D detections. MANO is a parametric model defined by shape parameters $\bm{\beta}\in\mathbb{R}^{10}$, pose parameters $\bm{\theta}\in\mathbb{R}^{45}$ (representing the axis-angle rotations of 15 joints), a global orientation $\mathbf{R}\in SO(3)$, and a global translation $\mathbf{t}\in\mathbb{R}^{3}$.
@@
$(\bm{\theta}^{*},\mathbf{R}^{*},\mathbf{t}^{*})=\arg\min_{\bm{\theta},\mathbf{R},\mathbf{t}}\frac{1}{N_{\text{cams}}}\sum_{c=1}^{N_{\text{cams}}}\frac{1}{21}\sum_{j=1}^{21}\left\|\Pi(\mathbf{K}_{c},[\mathbf{R}_{c}|\mathbf{t}_{c}],\mathbf{J}_{j}(\bm{\beta},\bm{\theta},\mathbf{R},\mathbf{t}))-\mathbf{k}_{c,j}\right\|_{2}^{2}$
@@
As detailed in our provided source code, this optimization problem is solved iteratively using the Adam optimizer. For any given subject, the shape parameter $\bm{\beta}$ is fixed to the mean shape ($\bm{\beta}=\mathbf{0}$) to ensure temporal consistency. We initialize the hand in a neutral pose and optimize the pose ($\bm{\theta}$), global orientation ($\mathbf{R}$), and translation ($\mathbf{t}$) parameters for each frame. The learning rate is managed by a Cosine Annealing scheduler, starting at 0.14 and decaying to 0.08. To balance accuracy and computational load, we employ a dynamic iteration scheduler that reduces the number of optimization steps for later frames in a sequence, assuming smaller inter-frame motion. This process yields a complete, kinematically consistent 3D representation of the hand pose. The full algorithm is summarized in Algorithm 1.
Algorithm 1 Model-Based 3D Hand Reconstruction
1:
3:
4:Optimized parameters $\bm{\theta}^{*},\mathbf{R}^{*},\mathbf{t}^{*}$.
5:
6:procedure ReconstructHandPose($\{\mathbf{k}_{c,j}\},\dots$)
7:  $\bm{\theta}\leftarrow\mathbf{0},\mathbf{R}\leftarrow\mathbf{I},\mathbf{t}\leftarrow\mathbf{0}$ $\triangleright$ $\triangleright$ Initialize optimizable parameters
8:  $\bm{\beta}\leftarrow\mathbf{0}$ $\triangleright$ $\triangleright$ Use fixed mean hand shape
9:  $N_{iter}\leftarrow\textsc{IterationScheduler.get\_iterations}$
10:  for $i=1\to N_{iter}$ do
11:   $\{\mathbf{J}_{j}\}\leftarrow\mathcal{M}(\bm{\beta},\bm{\theta},\mathbf{R},\mathbf{t})$ $\triangleright$ $\triangleright$ Generate 3D joints via MANO forward pass
13:   Update parameters $\bm{\theta},\mathbf{R},\mathbf{t}$ via gradient descent on lossloss.
14:  end for
15:  return $\bm{\theta},\mathbf{R},\mathbf{t}$ as $\bm{\theta}^{*},\mathbf{R}^{*},\mathbf{t}^{*}$
16:end procedure
@@
A.2.2 Robot Data Collection System
Hardware Setup. The robotic setup consists of a Franka Research 3 arm with a UMI Gripper  (Chi et al., 2024) as the end-effector. For intuitive control, we employ a 6-DoF SpaceMouse as the primary input device for the operator. The visual data acquisition system includes two synchronized RGB-D cameras. Including a static main camera positioned to the robot’s left (called Left Camera) and a wrist-mounted camera providing an ego-centric view (called Ego Camera). Notably, since it is easy to be occluded by the robot arm, the Top Camera is not used in the robot data collection system.
@@ A.3 Camera Calibration
A.4 Evaluation Metrics and Protocol
A.4.1 Evaluation Protocol Details
@@
A.4.2 Waypoint Definitions for Task Progress
The 8-step waypoint decomposition for the task progress metric is defined in Table 3.
@@ A.4.2 Waypoint Definitions for Task Progress
Table 3: Waypoint definitions for the two Long-Horizon tasks.
A.4.3 Evaluation System
@@ A.5 Model Training
We initialize our model using the pre-trained weights from $\pi_{0}$ (Black et al., ), a state-of-the-art vision-language-action model. The vision encoder is kept frozen during training to leverage its pre-learned visual representations. The training process employs the AdamW optimizer with a learning rate of $1\times 10^{-4}$, and a linear learning rate scheduler with warmup and decay phases. The parameters of trajectory expert ($g_{\mathcal{T}}$) is set to be the same as action expert ($\pi_{0}$), except for the output dimension.
The training framework is adapted from LeRobot (Cadene et al., 2024), with modifications to accommodate our multi-view input and dual expert architecture. We utilize a batch size of 32, with each training iteration processing chunks of 50 timesteps for both actions and trajectories.
The detailed hyperparameter configuration is presented in Table 4.
@@
Base Model Type $\pi_{0}$ (Black et al., )
@@
Table 4: Model Training Hyperparameter Details
A.6 Additional Visualization and Qualitative Analysis.
For the visualization of task pick up the tomato and put it in the tray and stack the rings on the pillar, please refer to Figure 4. For the visualization of task pick up the water bottle and stack the paper cups, please refer to Figure 9.

### v2

@@ IV Experiments
Data Collection Efficiency Analysis. As presented in Table I, a summary of our data collection effort reveals a significant disparity in collection efficiency between human hand and robot teleoperation. For each demonstration, we measure both Collect Time (the execution duration of the task itself) and Reset Time (the period needed to restore the robot, human and objects for the next trial). Overall, Collect Time from a human expert is roughly 3.5 times faster per demonstration than robot teleoperation (an average of 5.09s vs. 17.98s).
@@
TABLE I: Statistics of the collected demonstration data across four different tasks. We report the number of demonstrations (# Demos), time cost in minutes, and the data collection efficiency in demonstrations per second (#/s). Efficiency is reported for the collection phase, reset phase, and the total process.
@@
Task Progress (TP): For the more complex, long-horizon tasks (stack the rings on the pillar and stack the paper cups), a simple binary success metric can be too sparse. We therefore measure task progress to provide a finer-grained evaluation. Each task is decomposed into 8 key waypoints, and the policy earns one point for each waypoint it successfully completes. The final score is the average number of completed waypoints across all trials, normalized by the total of 8 waypoints. The detailed waypoint definitions are provided in Appendix A-C1.
Evaluation Protocol in Real-World Tasks. Each task was evaluated over 50 trials with varied initial object configurations (position, orientation) and the introduction of distractors to test for generalization.
IV-A Performance Analysis
@@
TABLE II: Performance comparison. Success Rate (%) for Short-Horizon (SH) tasks and Task Progress (%) for Long-Horizon (LH) tasks are reported. Baseline represents the pre-trained $\pi_{0}$ with task-specific finetuning.
Baseline. We establish a strong baseline using the $\pi_{0}$ architecture [6], initialized with its publicly available pre-trained weights. In line with standard VLA evaluation [7], we then fine-tune this model for each downstream task using our collected robot demonstrations. To ensure a fair and direct comparison, the amount of robot data used for fine-tuning the baseline is identical to the robot data subset used for training our full Traj2Action model. This setup provides a rigorous performance benchmark to precisely measure the benefits introduced by our trajectory-guided architecture and the integration of human data.
Contribution of Trajectory Expert. We conduct a two-part ablation study to dissect the contributions of trajectory expert. First, we analyze the impact of our architectural modification. To do this, we evaluate a version of our framework that includes the trajectory expert to baseline but is trained solely on robot data. As shown in Table II, simply adding the trajectory expert improves performance. The benefit is most pronounced on the long-horizon SPC task, where this architectural change alone boosts the Task Progress score from 37.75% to 54.25% (+16.50%). We attribute this to the explicit decomposition of the policy: the trajectory expert acts as a high-level planner by generating a coarse spatio-temporal plan. This plan serves as a robust prior that simplifies the problem for the action expert, allowing it to focus on refining the motion into precise, low-level actions.
@@
Contribution of Human Data. We then analyze the impact of incorporating human demonstrations for robot learning. As shown in Table II, integrating human data provides a further, substantial performance boost across all tasks. Compared to the baseline, our full Traj2Action model achieves significant gains: +28% on the pick up the water bottle task (reaching 76.00%), +26.00% on the pick up the tomato and put it in the tray task (reaching 76.00%), +23.50% on the stack the paper cups task (reaching 61.25%), and +21.00% on the stack the rings on the pillar task (reaching 44.75%). This demonstrates that long-horizon tasks, in particular, benefit from the diversity of human data. This is because human demonstrators can intuitively execute a wider range of motions—often involving agile re-orientations or angles that are awkward and difficult to perform via teleoperation. By supplementing the robot dataset with these more varied human demonstrations, the trajectory expert learns to predict more robust and generalizable plans. Consequently, the action expert’s performance is boosted. This effect is also qualitatively illustrated in Figure 4. In the long-horizon stack the rings on the pillar task, the baseline’s poor long-term plan causes it to get trapped. In contrast, our policy predicts a coherent trajectory and executes the task successfully. Furthermore, our policy demonstrates superior robustness: after an initial error in the pick up the tomato and put it in the tray task, it successfully re-plans and recovers, unlike the baseline which enters a non-productive loop.
Figure 4: Visual comparison of trajectory and action prediction of short- and long-horizon tasks pick up the tomato and put it in the tray (top) and stack the rings on the pillar (bottom), respectively.
IV-B Ablation Studies
@@
Strategy Robot Data (#/min) Human Data (#/min) Total Data Time (min) Performance SR(%)
Baseline 408/202.29 202.29 50
+ Trajectory Expert + Robot Data-Only 408/202.29 0/0 202.29 60
+ Trajectory Expert + Human & Robot Data 270/133.70 120/14.25 147.95 58
240/28.61 162.31 60
635/75.39 209.09 62
TABLE III: Performance comparison for the pick up the tomato and put it in the tray task under different data collection strategies. The Robot and Human Data columns show the number of demonstrations collected and the total required time.
@@
Human Data act as an Effective Substitute for Robot Data. To investigate whether labor-saving collected human data can replace time-consuming robot data without degrading policy training performance, we conduct an experiment with three configurations as presented in Table III.
The results show that human data can be leveraged to train robot policy efficiently. By leveraging just 240 human demonstrations and less robot data, our policy achieves an identical 60% success rate performance compared with model trained solely by robot data, but with a 20% data collection time reduction.
@@
Impact of Trajectory Sampling Frequency.
A critical challenge in unifying human and robot data is the inherent mismatch in their kinematic speeds; human movements are typically much faster than robot motions. We hypothesized that to sample trajectory data from human and robot with a consistent speed profile (i.e., where the spatial distance between consecutive points is similar), it is necessary to sample the faster human motion at a higher frequency than the slower robot motion.
@@ V Conclusion
VI Limitation
Our method exhibits a representative failure case, detailed in Appendix A-F, in which the policy successfully plans and executes a trajectory to the vicinity of the target object yet fails to complete the grasp despite multiple attempts. Though the predicted trajectory accurately guides the end-effector to the target spatial region, small residual errors in the contact phase preclude the successful execution of the grasp. This case demonstrates that the trajectory representation is sufficient for high-level motion planning but insufficient for precise contact-level interaction, highlighting the lingering challenges in fine-grained robotic control at object contact boundaries.
@@ Appendix A Appendix
A-A Data Collection System Details
@@
A-A1 Human Hand Motion Capture System
@@
Hardware Setup. The system is built around four synchronized cameras operating at 30 Hz. Three of these are static industrial cameras mounted on a rigid frame surrounding the data collection workspace, providing top-down, side-left, and side-right perspectives of the user’s hand. This arrangement is critical for minimizing self-occlusion. The fourth is a lightweight camera mounted on the back of the user’s hand, providing an ego-centric view analogous to the wrist camera on our robot.
@@
  • 2D Keypoint Detection: We leverage the robust Google MediaPipe [43] Hand Landmarker to process each video stream independently. For each hand detected in an image, the model outputs 21 2D landmarks ($\mathbf{u},\mathbf{v}$) in pixel coordinates. These landmarks serve as the 2D evidence for our 3D reconstruction.
  • Model-Based 3D Reconstruction: We fit the MANO [32] model to these multi-view 2D detections. MANO is a parametric model defined by shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$, pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{45}$ (representing the axis-angle rotations of 15 joints), a global orientation $\mathbf{R}\in SO(3)$, and a global translation $\mathbf{t}\in\mathbb{R}^{3}$.
@@
$\begin{split}(\boldsymbol{\theta}^{*},\mathbf{R}^{*},\mathbf{t}^{*})=\arg\min_{\boldsymbol{\theta},\mathbf{R},\mathbf{t}}&\frac{1}{N_{\text{cams}}}\sum_{c=1}^{N_{\text{cams}}}\frac{1}{21}\sum_{j=1}^{21}\bigg\|\Pi\big(\mathbf{K}_{c},[\mathbf{R}_{c}|\mathbf{t}_{c}],\\ &\mathbf{J}_{j}(\boldsymbol{\beta},\boldsymbol{\theta},\mathbf{R},\mathbf{t})\big)-\mathbf{k}_{c,j}\bigg\|_{2}^{2}\end{split}$ (3)
@@
As detailed in our provided source code, this optimization problem is solved iteratively using the Adam optimizer. For any given subject, the shape parameter $\boldsymbol{\beta}$ is fixed to the mean shape ($\boldsymbol{\beta}=\mathbf{0}$) to ensure temporal consistency. We initialize the hand in a neutral pose and optimize the pose ($\boldsymbol{\theta}$), global orientation ($\mathbf{R}$), and translation ($\mathbf{t}$) parameters for each frame. The learning rate is managed by a Cosine Annealing scheduler, starting at 0.14 and decaying to 0.08. To balance accuracy and computational load, we employ a dynamic iteration scheduler that reduces the number of optimization steps for later frames in a sequence, assuming smaller inter-frame motion. This process yields a complete, kinematically consistent 3D representation of the hand pose. The full algorithm is summarized in Algorithm 1.
@@
A-A2 Robot Data Collection System
Hardware Setup. The robotic setup consists of a Franka Research 3 arm with a UMI Gripper  [10] as the end-effector. For intuitive control, we employ a 6-DoF SpaceMouse as the primary input device for the operator. The visual data acquisition system includes two synchronized RGB-D cameras. Including a static main camera positioned to the robot’s left (called Left Camera) and a wrist-mounted camera providing an ego-centric view (called Ego Camera). Notably, since it is easy to be occluded by the robot arm, the Top Camera is not used in the robot data collection system.
@@ A-B Camera Calibration
A-C Evaluation Metrics and Protocol
A-C1 Evaluation Protocol Details
@@
A-C2 Waypoint Definitions for Task Progress
The 8-step waypoint decomposition for the task progress metric is defined in Table IV.
@@ A-C2 Waypoint Definitions for Task Progress
TABLE IV: Waypoint definitions for the two Long-Horizon tasks.
A-C3 Evaluation System
@@ A-D Model Training
We initialize our model using the pre-trained weights from $\pi_{0}$ [6], a state-of-the-art vision-language-action model. The vision encoder is kept frozen during training to leverage its pre-learned visual representations. The training process employs the AdamW optimizer with a learning rate of $1\times 10^{-4}$, and a linear learning rate scheduler with warmup and decay phases. The parameters of trajectory expert ($g_{\mathcal{T}}$) is set to be the same as action expert ($\pi_{0}$), except for the output dimension.
The training framework is adapted from LeRobot [7], with modifications to accommodate our multi-view input and dual expert architecture. We utilize a batch size of 32, with each training iteration processing chunks of 50 timesteps for both actions and trajectories.
The detailed hyperparameter configuration is presented in Table V.
@@
Base Model Type $\pi_{0}$ [6]
@@
TABLE V: Model Training Hyperparameter Details
A-E Additional Visualization and Qualitative Analysis.
For the visualization of task pick up the water bottle and stack the paper cups, please refer to Figure 8.
@@
A-F Limitation
Figure 9 illustrates some representative failure cases of our method on the task pick up the tomato and put it in the blue tray.
Figure 9: Representative failure cases on the task pick up the tomato and put it in the blue tray.